OpenStudy (anonymous):

multiply the denominators (1+√5)(1-√5)

OpenStudy (anonymous):

the conjugates and simplify ???

OpenStudy (anonymous):

??? Denominators? Conjugates? \[(1+\sqrt{5})(1-\sqrt{5}) = -4\]

OpenStudy (gw2011):

Actually, (1+square root 5) and (1-square root 5) are conjugates of each other. Since you mention denominators, do you mean the problem should be as follows: (1+square root 5)/(1-square root 5)? If so, then the following solution can be arrived at: [(1+square root 5)/(1-square root 5)]*[(1+square root 5)/(1+square root 5)] [1+2(square root 5)+5]/(1-5)=[6+2(square root 5]/-4 Therefore, the answer is: -(3+square root 5)/2